OpenStudy (anonymous):

let f(x)=-5x-4 and give g(x)=6x-7. find f(x)=g(x)

OpenStudy (jtug6):

They want when both equations equal eachother so whats -5x - 4 = 6x-7 where we are solving x?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

these are the answer choices -11x+3 x+3 -11x-11 x-11

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ f(x) }} =-5x-4 \qquad {\color{blue}{ g(x) }} =6x-7 \\ \quad \\ f(x)=g(x)\implies {\color{brown}{ -5x-4 }}={\color{blue}{ 6x-7 }}\implies x=?\) so.. what does that give you?

OpenStudy (anonymous):

i dont know, i dont know how to do this

OpenStudy (jdoe0001):

well... how come you have been given the exercise, without covering the subject first then?

OpenStudy (jtug6):

^

OpenStudy (anonymous):

i came here for help not to be criticized